OpenStudy (anonymous):

Someone help? A function that can be written in the form f(x)=ax^2+bx+c , where a does not=0 is called a __________ function. Its graph is a shape called a __________.

OpenStudy (anonymous):

quadratic; parabola

OpenStudy (anonymous):

Even if a doesn't equal 0?

OpenStudy (anonymous):

It says where a is not 0.

OpenStudy (anonymous):

as long as there is an x^2 term, it is quadratic. The following are all quadratic equations: x^2 - 3x + 9 = 0 3x^2 - 8 = 0 2x^2 = 4x

OpenStudy (anonymous):

Oh ok thanks for explaining, im working an a really long mid-term review right now